OpenStudy (anonymous):

a ball is hit into the air; its height, in metres. after t seconds is h = -5^2 + 30t + 1 for how long is the ball in the air? how do i find the x intercepts when it cannot be factored?

OpenStudy (sumeet):

somethings missing in this. is the ball hit at inclined angle , something like projectile motion. Then it should have t^2 term in it, the path being parabolic. you should get two values of t for each height, one is the time when it reaches height while going up and the other time while going down through the same height

OpenStudy (agreene):

I'm assuming you meant: \[h(t)= -5t^2 + 30t + 1 \] \[0= -5t^2+30t+1\] \[0=t^2-6t-\frac15\] \[\frac{46}5 = t^2 - 6t +9\] \[\frac{46}5 = (t-3)^2\] \[t= \frac15(15\pm\sqrt{230})\] \[t\approx-0.0331; \space\space\space t\approx 6.0331\]